1. Suppose you are an event coordinator for a large performance theater. One of the hottest new Broadway musicals has started to tour, and your city is the first stop on the tour. You need to supply information about projected ticket sales to the box office manager. The box office manager uses this information to anticipate staffing needs until the tickets sell out. You provide the manager with a quadratic equation that models the expected number of ticket sales for each day x. ( is the day tickets go on sale).

a. Does the graph of this equation open up or down? How did you determine this?

b. Describe what happens to the tickets sales as time passes?
c. Use the quadratic equation to determine the last day that tickets will be sold. (Note: Write your answer in terms of the number of days after ticket sales begin.)

d. Will tickets peak or be at a low during the middle of the sale? How do you know?

e. After how many days will the peak or low occur?

f. How many tickets will be sold on the day when the peak or low occurs?

g. What is the point of the vertex? How does this number relate to your answers in parts e and f?

h. How many solutions are there to the equation ? How do you know?

i. What do the solutions represent? Is there a solution that does not make sense? If so, in what ways does the solution not make sense?

Answer by stanbon(75887)

a. Does the graph of this equation open up or down? How did you determine this?
My Opinion: Ticket sales rise to a peak then fall.
The graph opens down.
b. Describe what happens to the tickets sales as time passes?
Ticket sales rise to a peak then fall.

c. Use the quadratic equation to determine the last day that tickets will be sold. (Note: Write your answer in terms of the number of days after ticket sales begin.)
Sales(x) = -x(x-23)
That says "zero sales on day zero and zero sales on day 23.
----------------------
d. Will tickets peak or be at a low during the middle of the sale? How do you know?
Peak. It's a matter of opinion.
e. After how many days will the peak or low occur?
The 11th day.
f. How many tickets will be sold on the day when the peak or low occurs?
S(11) = -11(11-23) = -11*-12 = 132 tickets sold at the peak; zero at the low.

g. What is the point of the vertex? How does this number relate to your answers in parts e and f?
(11,132)
h. How many solutions are there to the equation ? How do you know?
Two: x = 0 or x = 23
i. What do the solutions represent? Is there a solution that does not make sense? If so, in what ways does the solution not make sense?
x>23 makes no sense; ticket sales will be negative.
======================================
Cheers,
Stan H.